# Tripartite Entanglement In Mixed-Spin Triangle Trimer

### v1

@@ 4 Results
According to the definition given above (Eq. (7)) we have obtained analytic expressions for the tripartite negativity for the eingenstates of the system under consideration. We present here only those which are relevant for the ground states phase diagram, obtained in our previous work [40]. As tripartite negativity is given by the Eq. (7) it is enough to know only all generalized bipartite negativities, $Ne^{i}_{A_{i}-A_{j}A_{k}}$. The index $i$ stands for the eigenstate numbering parameter. Bellow, we denote spin-1/2 ion with $g$-factor equal to $g_{1}$ by $A$, spin-1/2 ion with $g$-factor equal to $g_{2}$ by $B$ and ion with spin-1 by $C$. Interestingly, some eigenstates can have the same $Ne^{i}_{A_{i}-A_{j}A_{k}}$.
$Ne^{3,4}_{A-BC}=Ne^{3,4}_{C-AB}=\frac{\sqrt{2}}{3},\>\>\>\>\>\>Ne^{3,4}_{B-AC}=0,\>\>\>\>\>\>Ne^{5,6}_{A-BC}=\frac{\sqrt{2M_{\pm}^{4}+5M_{\pm}^{2}+2}}{3\left(M_{\pm}^{2}+1\right)},$ (11)
@@
$Ne^{11,12}_{B-AC}=\frac{\sqrt{U_{\mp}^{2}+2}}{U_{\mp}^{2}+3},\>\>\>\>\>\>Ne^{11}_{C-AB}=\frac{\sqrt{2}\sqrt{U_{\mp}^{2}+1}}{U_{\mp}^{2}+3}.$
In our previous work a detailed analysis of the ground states phase diagrams of the system were presented [40]. Here we demonstrate only those of them, which exhibit interesting features of the tripartite negativity dependent on the value on non-uniform $g$-factor. The main purpose of the present research is to figure out how non-uniform $g$-factors affect tripartite entanglement properties of the model. It is worth mentioning, that the most remarkable enhancement of the bipartite negativity caused by non-uniform $g$-factors reported in our previous work [40] concerned the case $J_{2}=J_{1}$ and gave almost 7-fold robust decrease of $Ne_{12}$ for arbitrary small difference between $g_{2}$ and $g_{1}$. For the tripartite entanglement situation is quite different. Here we presented density plots of the tripartite negativity, $Ne_{ABC}$, projected onto ground state phase diagrams in the ”$g$-factors ration” - ”dimensionless magnetic field” plane.
Figure 2: Density plots of tripartite negativity for $J_{1}=J_{2}>0$ (panel a), $J_{2}>0\,\,\text{and}\,\,J_{1}=\frac{1}{5}J_{2}$ (panel b) and $J_{2}>0\,\,\text{and}\,\,J_{1}=-\frac{1}{5}J_{2}$ (panel c) cases.
@@
Two antiferromagnetic cases, $J_{1}=J_{2}>0$, $J_{2}>0\,\,\text{and}\,\,J_{1}=\frac{1}{5}J_{2}$, and a mixed case, $J_{2}>0\,\,\text{and}\,\,J_{1}=-\frac{1}{5}J_{2}$, are presented in the Fig. 2. In the case of equal antiferromagnetic coupling (panel a) the phase diagram includes four eigenstates $|\psi_{1}\rangle,|\psi_{3}\rangle,|\psi_{5}\rangle$ and $|\psi_{9}\rangle$. However, two of them $|\psi_{3}\rangle$ and $|\psi_{9}\rangle$ transform non-continuously under $g_{2}\rightarrow g_{1}$. For uniform situation, $g_{2}=g_{1}$, degeneracy line between $|\psi_{3}\rangle$ and $|\psi_{9}\rangle$ regions corresponds to the degenerate superposition of $|\psi_{3}\rangle_{0}$ and $|\psi_{9}\rangle_{0}$ given in the Eq. (4). This segment is highlighted in green in the right panel on the Fig. 2. The value of the tripartite negativity here is $Ne_{ABC}^{(3+9)_{0}}=\frac{1}{2}\sqrt[3]{\frac{3}{4}}\approx 0.45$. Interestingly, the $|\psi_{5}\rangle$ eigenstate, which is the zero-field ground state for arbitrary value of $g_{2}/g_{1}$, exhibits maximal three-particle entanglement at $B=0$ and at the segment $g_{2}=g_{1}$, $Ne_{ABC}^{5}=\sqrt[3]{\frac{1}{4}}\approx 0.63$. The most dramatic discrepancy between bipartite and tripartite entanglement properties is occurred for the $|\psi_{3}\rangle$ eigenstate, for which one of the bipartite negativities becomes almost seven times larger than the corresponding uniform-$g$ value [40]. Tripartite negativity is zero for $|\psi_{3}\rangle$, which means that this state is biseparable. Thus, in the case of $g_{2}>g_{1}$, when magnetization is not conserved, it is possible to create two significantly different entanglement regimes and control them by external magnetic field. First regime exhibits maximum value for tripartite entanglement, when the second regime exhibits maximum value for bipartite entanglement and have no tripartite entanglement. As usual, for the large enough magnetic field the system reaches its saturated ($|\psi_{1}\rangle$) or quasi-saturates ($|\psi_{9}\rangle$) states with zero or vanishing entanglement. For the case $J_{2}>0,\,\,J_{1}=\frac{1}{5}J_{2}$ (Fig. 2, panel b) system exhibits similar behavior, there are eigenstates with strong bipartite entanglement, but with zero $Ne_{ABC}$, $|\psi_{3}\rangle$ and region where both quantities are quite large, $|\psi_{5}\rangle$. However, presence of non-conserving magnetization does not bring any quantitative or qualitative benefits. The only difference between $g_{2}=g_{1}$ and $g_{2}\neq g_{1}$ cases are that for the second case tripartite negativity is depended on magnetic field within the same ground state. Fig. 2 (panel c) shows density plot of tripartite negativity for the $J_{2}>0,\,\,J_{1}=-\frac{1}{5}J_{2}$ case. Here system have one additional region of ground state corresponding to $|\psi_{4}\rangle$. Here in contrast to the previous cases, regime with maximal or essentially large $Ne_{ABC}$ is absent. System exhibits maximal entanglement under the conditions $g_{2}=-g_{1}$ and when magnetic field is close to zero. In the Fig. 3 the cases $J_{1}<0$, arbitrary $J_{2}\leq 0$ and $J_{1}>0$, arbitrary $J_{2}<0$ are presented. For both cases maximal possible tripartite negativity achieved for the eigenstate $|\psi_{7}\rangle$ with value 0.59.
Figure 3: Density plots of tripartite negativity for $J_{1}<0$, arbitrary $J_{2}\leq 0$ and $J_{1}>0$, arbitrary $J_{2}<0$ cases in panels a and b respectfully.
@@ 5 Conclusion
In the paper we considered the mixed spin (1/2, 1/2, 1) Heisenberg spin trimer with non-conserving magnetization. Analytical results for tripartite negativity were obtained for all ground states. In general, depending on the relations between two exchange constants, the system can exhibit five regimes of magnetic behavior [40]. These regimes, in their turn, are storngly affected by the ration of the $g$-factors, $g_{2}/g_{1}$. It was shown that in fully antiferromagnetic cases, it is possible to obtain a three-regime system (fully separable, biseparable, tripartite entangled) and control it by magnetic field. In the case $J_{2}=J_{1}>0$ these regimes are possible only when $g_{2}>g_{1}$, when magnetisation is not-conserving.
6 Acknowledgements
The authors acknowledge partial financial support form ANSEF (Grants No. PS-condmatth-2462 and PS-condmatth-2884) and from CS RA MESCS (Grants No. 21AG-1C047, 21AG-1C006 and 23AA-1C032).

### v2

@@ 4 Results
According to the definition given above [equation (3.3)] we have obtained analytic expressions for the tripartite negativity for the eingenstates of the system under consideration. We present here only those which are relevant for the ground states phase diagram, obtained in our previous work [39]. Since tripartite negativity is given by the equation (3.3), it is sufficient to know only all generalized bipartite negativities, $Ne^{i}_{A_{i}-A_{j}A_{k}}$. The index $i$ stands for the eigenstate numbering the parameter. Bellow, we denote spin-1/2 ion with $g$-factor equal to $g_{1}$ by $A$, spin-1/2 ion with $g$-factor equal to $g_{2}$ by $B$ and ion with spin-1 by $C$. Interestingly, some eigenstates can have the same $Ne^{i}_{A_{i}-A_{j}A_{k}}$.
Figure 2: (Colour online) Density plots of tripartite negativity for $J_{1}=J_{2}>0$ (panel a), $J_{2}>0\,\,\text{and}\,\,J_{1}=\frac{1}{5}J_{2}$ (panel b) and $J_{2}>0\,\,\text{and}\,\,J_{1}=-\frac{1}{5}J_{2}$ (panel c) cases.
Figure 3: (Colour online) Density plots of tripartite negativity for $J_{1}<0$, arbitrary $J_{2}\leqslant 0$ and $J_{1}>0$, arbitrary $J_{2}<0$ cases in panels a and b, respectively.
$Ne^{3,4}_{A-BC}=Ne^{3,4}_{C-AB}=\frac{\sqrt{2}}{3},\>\>\>\>\>\>Ne^{3,4}_{B-AC}=0,\>\>\>\>\>\>Ne^{5,6}_{A-BC}=\frac{\sqrt{2M_{\pm}^{4}+5M_{\pm}^{2}+2}}{3\left(M_{\pm}^{2}+1\right)},$
@@
$Ne^{11,12}_{B-AC}=\frac{\sqrt{U_{\mp}^{2}+2}}{U_{\mp}^{2}+3},\>\>\>\>\>\>Ne^{11,12}_{C-AB}=\frac{\sqrt{2}\sqrt{U_{\mp}^{2}+1}}{U_{\mp}^{2}+3}.$ (4.1)
@@
In our previous work, a detailed analysis of the ground states phase diagrams of the system were presented [39]. Here, we demonstrate only those of them, which exhibit interesting features of the tripartite negativity dependent on the value on non-uniform $g$-factor. The main purpose of the present research is to figure out how non-uniform $g$-factors affect tripartite entanglement properties of the model. It is worth mentioning that the most remarkable enhancement of the bipartite negativity caused by non-uniform $g$-factors reported in our previous work [39] concerned the case $J_{2}=J_{1}$ and gave almost 7-fold robust increase of $\mathit{Ne}_{13}$ for arbitraryly small difference between $g_{2}$ and $g_{1}$. For the tripartite entanglement, the situation is quite different. Here, we presented density plots of the tripartite negativity, $\mathit{Ne}_{ABC}$, projected onto ground state phase diagrams in the “$g$-factors ration” – “dimensionless magnetic field” plane. Two antiferromagnetic cases, $J_{1}=J_{2}>0$, $J_{2}>0\,\,\text{and}\,\,J_{1}=\frac{1}{5}J_{2}$, and a mixed case, $J_{2}>0\,\,\text{and}\,\,J_{1}=-\frac{1}{5}J_{2}$, are presented in the figure 2. In the case of equal antiferromagnetic coupling (panel a) the phase diagram includes four eigenstates $|\psi_{1}\rangle,|\psi_{3}\rangle,|\psi_{5}\rangle$ and $|\psi_{9}\rangle$. However, two of them, $|\psi_{3}\rangle$ and $|\psi_{9}\rangle$, transform non-continuously under $g_{2}\rightarrow g_{1}$. For uniform situation, $g_{2}=g_{1}$, degeneracy line between $|\psi_{3}\rangle$ and $|\psi_{9}\rangle$ regions corresponds to the degenerate superposition of $|\psi_{3}\rangle_{0}$ and $|\psi_{9}\rangle_{0}$ given in the equation (2.4). This segment is highlighted in green in the right-hand panel in figure 2. The value of the tripartite negativity here is $\mathit{Ne}_{ABC}^{(3+9)_{0}}=\frac{1}{2}\sqrt[3]{\frac{3}{4}}\approx 0.45$. Interestingly, the $|\psi_{5}\rangle$ eigenstate, which is the zero-field ground state for arbitrary value of $g_{2}/g_{1}$, exhibits maximal three-particle entanglement at $B=0$ and at the segment $g_{2}=g_{1}$, $\mathit{Ne}_{ABC}^{5}=\sqrt[3]{\frac{1}{4}}\approx 0.63$. The most dramatic discrepancy between bipartite and tripartite entanglement properties occurrs for the $|\psi_{3}\rangle$ eigenstate, for which one of the bipartite negativities becomes almost seven times larger than the corresponding $g$-uniform value [39]. Tripartite negativity is zero for $|\psi_{3}\rangle$, which means that this state is biseparable. Thus, in the case of $g_{2}>g_{1}$, when magnetization is not conserved, it is possible to create two considerably different entanglement regimes and control them by an external magnetic field. The first regime exhibits maximum value for tripartite entanglement, while the second regime exhibits maximum value for bipartite entanglement and has no tripartite entanglement. As usual, for a large enough magnetic field, the system reaches its saturated ($|\psi_{1}\rangle$) or quasi-saturated ($|\psi_{9}\rangle$) states with zero or vanishing entanglement. For the case $J_{2}>0,\,\,J_{1}=\frac{1}{5}J_{2}$ (figure 2, panel b) the system exhibits a similar behavior. There are eigenstates with strong bipartite entanglement, but with zero $\mathit{Ne}_{ABC}$, $|\psi_{3}\rangle$ and the region where both quantities are quite large, $|\psi_{5}\rangle$. However, the presence of non-conserving magnetization does not bring any quantitative or qualitative benefits. The only difference between $g_{2}=g_{1}$ and $g_{2}\neq g_{1}$ cases are that for the second case tripartite negativity is dependent on the magnetic field within the same ground state. Figure 2 (panel c) shows the density plot of tripartite negativity for the $J_{2}>0,\,\,J_{1}=-\frac{1}{5}J_{2}$ case. Here, the system has one additional region of ground state corresponding to $|\psi_{4}\rangle$. Here, in contrast to the previous cases, the regime with maximal or essentially large $\mathit{Ne}_{ABC}$ is absent. The system exhibits maximal entanglement under the conditions $g_{2}=-g_{1}$ and when magnetic field is close to zero. In the figure 3 the cases $J_{1}<0$, arbitrary $J_{2}\leqslant 0$ and $J_{1}>0$, arbitrary $J_{2}<0$ are presented. For both cases, maximal possible tripartite negativity is achieved for the eigenstate $|\psi_{7}\rangle$ with value 0.59.
@@ 5 Conclusion
In the paper we considered the mixed spin (1/2, 1/2, 1) Heisenberg spin trimmer with non-conserving magnetization. Analytical results for tripartite negativity were obtained for all ground states. In general, depending on the relations between two exchange constants, the system can exhibit five regimes of magnetic behavior [39]. These regimes, in their turn, are storngly affected by the ration of the $g$-factors, $g_{2}/g_{1}$. It was shown that in fully antiferromagnetic cases, it is possible to obtain a three-regime system (fully separable, biseparable, tripartite entangled) and control it by magnetic field. In the case $J_{2}=J_{1}>0$, these regimes are possible only when $g_{2}>g_{1}$, when magnetisation is non-conserving.
Acknowledgements
The authors acknowledge partial financial support form ANSEF (Grants No. PS-condmatth-2884 and PS-condmatth-3273 ) and from CS RA MESCS (Grants No. 21AG-1C047, 21AG-1C006 and 23AA-1C032).